OpenStudy (howard-wolowitz):

If you help me I'll medal :

OpenStudy (anonymous):

I will help you get started 3 log(2x) = 4 log (2x)^3 = 4 raise both sides to the power of 10 (2x)^3 = 10^4 (2x) = (10^4)^(1/3)

OpenStudy (mrnood):

Same answer - different sequence: First divide both sides by 3 log 2x = 4/3 10 to power both sides: 2x = 10^ (4/3)